OpenStudy (anonymous):

Calculate the derivative: ddivdx int_{x ^{2}}^{x ^{3}}sqrt{x}dt

OpenStudy (anonymous):

THIS IS THE QUESTION: \[ddivdx \int\limits_{x ^{2}}^{x ^{3}}\sqrt{x}dt\]

OpenStudy (anonymous):

it asks to fine the derivative

OpenStudy (anonymous):

**find**

OpenStudy (ash2326):

i'm not able to understand the question, could you post it again

OpenStudy (ash2326):

i get it\[d/dx(\int\limits_{x^2}^{x^3}\sqrt(t)dt)\]

OpenStudy (anonymous):

so integrate first then take the derivative with respect to the new x values you get from the limits of integration

OpenStudy (amistre64):

the derivative of an integration is simply the inegral argument itself. \[f(x)=\int_{a(x)}^{b(x)}f'(t)dt=f[b(x)]-f[a(x)]\] \[f'(x)=f[b(x)]'-f[a(x)]'=f'[b(x)]b'(x)-f'[a(x)]a'(x)\]

OpenStudy (amistre64):

the shortcut it to cut out the middle part since you get right back to the argument in the integration to begin with

OpenStudy (amistre64):

\[\sqrt{t}\text{ is your arguement}\] \[\sqrt{x^3}\ 3x^2-\sqrt{x^2}\ 2x\]